# Discovery of Seven Cold and Distant Brown Dwarfs with JWST RUBIES

### v1

@@ 3.2 Markov Chain Monte Carlo Fitting
We used the ucdmcmc package v1.1.2 (Burgasser et al., 2025) to fit the three model sets to our brown dwarf spectra. Spectra and models were compared using the same chi-square statistic as Eqn. 1. A first fit was conducted to identify the best individual model within a model set based on the minimum value. These model parameters were then used to initiate a Metropolis-Hastings Markov Chain Monte Carlo (MCMC) algorithm (Metropolis et al., 1953; Hastings, 1970) to determine the overall best-fit parameters and their uncertainties from the posterior distributions. Model logarithmic flux densities were linearly interpolated between grid points over logarithmic parameter values (e.g., interpolates as log ). The MCMC algorithm proposes a new set of parameters by varying the previous set using normal distributions with the fixed widths listed in Table 3. The difference between the values of the proposed () and previous () parameters were compared to the overall minimum , which in turn was compared to a random number drawn from a uniform distribution between 0 and a threshold value :
@@ 4 Results
Best fit models for each source and model set are displayed in Figures 4 and 5, and fit parameters are listed in Tables 4 and 5, for the 0.9–2.4 m (NIR) and 0.9–5.1 m (NIRSpec) regions, respectively. A summary of fit parameters across models is provided in Table 6, while a full list of parameter values and distribution visualizations are provided in a separate Zenodo link.111[placeholder URL for final version]. Here, we describe the outcomes of the fits and the physical interpretation of the fit parameters for individual sources.
@@ 4.1 RUBIES-BD-1
The best NIR and NIRSpec fits for this T7 dwarf are from the Elf Owl and Diamondback models, with a slight preference for the former with modest vertical mixing ( = 3.11.0 averaged between NIR and NIRSpec fits). The Diamondback models require very efficient sedimentation or no clouds ( = 9.50.4). Averaging the results of the NIRSpec fits for these models yields Teff = 930 K, = 4.90.4, and [M/H] = +0.1, while the Elf Owl models indicate a solar C/O ratio. The inferred effective temperature is slightly higher but consistent with previously established trends in spectral type (e.g., Teff = 830130 K for a T70.5 dwarf based on Filippazzo et al. 2015). This modest overshoot may be due to the best-fitting models failing to match the depth of the 1.6 m CH4 band or the strong CO2 band at 4.1 m (see discussion below). Nevertheless, the agreement in classification and model fit parameters indicate that this is a normal late T-type brown dwarf.
RUBIES-BD-1 was previously reported as CEERS-EGS-BD-4 in Hainline et al. (2024a), identified from multi-band JWST/NIRCam photometry. The photometric measurements were fit to three sets of atmosphere models: Sonora Bobcat and Sonora Cholla (Karalidi et al., 2021), precursors to Diamondback and Elf Owl models; and ATMO (Phillips et al., 2020). The fits yield Teffs spanning 1000–1050 K and = 5.0, but no evaluation of metallicity is reported. The authors infer a distance of 1350–1831 pc for estimated radii of 0.92–1.2 RJup, somewhat further than our estimate of 1030140 pc using a smaller radius estimate. RUBIES JWST/NIRSpec spectral data of this source were previously reported in Tu et al. (2025a), where the source is identified as o006_s00089. These authors compared the NIRSpec spectrum to Elf Owl and ATMO 2020++ (Meisner et al., 2023) model sets, the latter an updated version of the ATMO models. For the former, the authors inferred a similar Teff = 955 K but a considerably smaller = 4.34, outside the range considered in these fits (albeit formally consistent with our uncertainties). They also infer a much larger distance for this source, 2.00.2 kpc, by assuming a radius of 1.2 RJup as suggested by the low surface gravity, but estimate a shorter distance of 1.4–1.7 kpc if the radius is constrained to 0.8–0.9 RJup. The authors find a significantly supersolar metallicity of [M/H] = +0.410.05 from their model fits, above but again formally consistent with our fits. Tu et al. (2025a) note that their parameter uncertainties, which are considerably smaller than those reported here, may be underestimated.
@@ 4.2 RUBIES-BD-2
The NIR spectrum for this L1 dwarf is clearly best fit by the Diamondback models, while the full NIRSpec spectrum has equal values between the Diamondback and SAND models. As both models include a prescription for cloud condensation and scattering, it is likely the presence of clouds in the atmosphere of this source that rules out the condensate-free Elf Owl models as a reliable fit. Indeed, Diamondback model fits to both the NIR and NIRSpec spectra require relatively thick clouds ( = 1.81.0). However, despite providing the best fits, Figure 5 shows that the Diamondback and SAND models have difficulty reproducing the 0.9–2.4 m spectral slope of RUBIES-BD-2, the former underestimating the 1.6 and 2.1 m flux peaks and the latter overestimating the 1.2 and 1.6 m peaks. We also find that the fits between the NIR and NIRSpec ranges yield marginally distinct temperatures, with Teff = 1760 K for the former (Diamondback only) and Teff = 2040 K for the latter (average of Diamondback and SAND), a 1.7 difference. The higher temperature is more in line with local L10.5 dwarfs (Teff = 2100130 K from Filippazzo et al. 2015). We also find a 3 difference in the inferred value between the SAND models ( = 4.22) and Diamondback models ( = 5.30) based on fits to the NIRSpec data; and a 4 discrepancy between our model-scaled distance estimate of 860 pc and our spectrophotometric distance estimate of 1320100 pc. These deviations suggest that the Elf Owl, Diamondback, and SAND models all have difficulty reproducing the spectrum of RUBIES-BD-2, indicating potentially missing elements to the models.
RUBIES JWST/NIRSpec spectral data of this source were previously reported in Tu et al. (2025a), where the source is identified as o006_s35616. The authors compared these data to Elf Owl and BT-Settl CIFIST models (Allard et al., 2012; Caffau et al., 2011), finding Teff = 1810–1910 K, lower but consistent with our NIRSpec fits. However, the authors again infer a very low = 3.250.01 and a supersolar [M/H] = +1.000.01 from their Elf Owl fits, inconsistent with our findings. Tu et al. (2025a) note significant residuals in their model fits, recognizing that neither the Elf Owl nor BT-Settl models include condensate opacity. They introduce a reddening parameter based on the interstellar extinction law of Wang & Chen (2024), and find that = 3 mag substantially improves their Elf Owl model fits while yielding a substantially higher Teff = 2286 K and subsolar [M/H] = 0.710.02. It is likely that a combination of atmospheric condensate opacity and interstellar reddening are required to accurately reproduce these data.
@@ 4.3 RUBIES-BD-3
As discussed in Section 2, RUBIES-BD-3 is better matched to subdwarf standards, suggesting that it should be well-matched to metal-poor atmosphere models. However, both NIR and NIRSpec fits yield consistently better fits to supersolar metallicity models. In the NIR range, Elf Owl, Diamondback, and SAND models all yield comparable best-fit values; in the NIRSpec range, the Elf Owl models are somewhat less reliable. Primary fit parameters of Teff, and [M/H] are in agreement across all three models and fit ranges, with a relatively high Teff = 2170130 K from the NIRSpec fits, notably warmer than the earlier-type L dwarf RUBIES-BD-2. Secondary parameters are similar to NIRSpec fits for RUBIES-BD-2, with all three models indicating relatively low surface gravities ( = 4.70.4), and for the Diamondback models thick clouds ( = 2.1). The Elf Owl model fits also indicate a high C/O = 0.73 and modest vertical mixing ( = 3.6). However, like RUBIES-BD-2, all three models have clear flaws in reproducing the observed spectrum, particularly in the NIR range where Elf Owl models show excess flux shortward of 1.3 m, and SAND models fail to match the shape of the 1.6 m peak. None of the models accurately reproduce the 1.2 m “notch” absorption feature that is characteristic of L- and T-type subdwarfs (Figure 3). Again, the models deployed for this seemingly metal-poor L dwarf may be missing a critical element of atmospheric chemistry, composition, or dynamics that may be skewing the fits toward higher metallicities.
Tu et al. (2025a) previously reported RUBIES JWST/NIRSpec spectral data for this source, identified it as o005_s41280, and compared the data to Elf Owl and BT-Settl models. The authors find a similar Teff = 2160–2260 K from these fits without taking into account reddening, with an approximately 200 K increase when a modest reddening of = 1.0-1.5 mag is taken into account. In both cases, the authors find a very low surface gravity for this source, = 3.3–3.5, well outside our fitting parameter range. As with RUBIES-BD-2, such a low surface gravity is highly unlikely for a field brown dwarf as it would require an age 1 Myr (Chabrier et al., 2023).
@@ 4.4 RUBIES-BD-4
The relatively low S/N of the RUBIES-BD-4 spectrum results in equivalent matches among all three models in both NIR and NIRSpec ranges, and relatively large uncertainties in inferred parameters. The poorly-constrained Teff = 1800 K is higher but not inconsistent with trends among solar metallicity dwarfs for its T23 classification (Teff = 1180200 K; Filippazzo et al. 2015). Notably, SAND model fits to the NIRSpec data yield a considerably lower Teff = 1050 K as compared to the other models, even though the best-fit model has a Teff = 1506 K (Figure 5). Secondary parameters indicate efficient vertical mixing ( = 7.81.0 from Elf Owl models), thin clouds ( = 8.40.7 from Diamondback models) and near-solar elemental abundances ([M/H] = 0.1 from all models, C/O = 0.54 from Elf Owl models, and [/M] = +0.17 from SAND models). The large range in temperatures across the models results in a comparably large uncertainty in the estimated distance, from 2000 pc for SAND to 6400 pc for Diamondback, which straddle the similarly uncertain 5400 pc distance estimate inferred from its spectral type and NIRCam photometry (Table 2).
@@ 4.5 RUBIES-BD-5
RUBIES-BD-5 is best fit in both the NIR and NIRspec ranges by the SAND models, yielding an atmospheric temperature Teff = 1270 K that is in line with its T12 classification (Teff = 1220150 K; Filippazzo et al. 2015). Figure 5 shows that the SAND models do a better job reproducing the 1–2.2 m and 4.0 m peaks compared to the Elf Owl and Diamondback models, both of which produce overly blue spectra. The NIRSpec SAND fits also yield a significantly subsolar metallicity ([M/H] = 0.7) and modest alpha enrichment ([/M] = +0.19), both indicative of a thick disk or halo origin.
@@ 4.6 RUBIES-BD-6
RUBIES-BD-6, like RUBIES-BD-4, has a relatively low S/N spectrum and is equally-well fit by all three models in both NIR and NIRSpec ranges, with a slight preference for the Elf Owl models in the latter. The NIRSpec fits yield an average Teff = 830130 K, high but consistent with its T81 classification (Teff = 680190 K; Filippazzo et al. 2015). Given the large uncertainties, there is no evidence of a metallicity or elemental compositions significantly distinct from solar, while the Diamondback models require efficient sedimentation ( = 8.0) as is typical for late T dwarfs.
@@ 4.7 RUBIES-BD-7
RUBIES-BD-7 is best fit in both the NIR and NIRSpec ranges by the Elf Owl and Diamondback models, the latter requiring very efficient sedimentation ( = 9.60.4, consistent with no clouds). The inferred Teff = 118080 K is consistent with other T50.5 dwarfs (Teff = 1030120 K; Filippazzo et al. 2015), while the metallicity ([M/H] = 0.20.3) and Elf Owl C/O = 0.48 are consistent with solar composition. The excellent agreement between the model scaling and spectrophotometric distances further indicate that this is likely a normal thin disk brown dwarf.
@@
Model labels are: E = Elf Owl (Mukherjee et al., 2024), D = Diamondback (Morley et al., 2024), and S = SAND (Gerasimov et al., 2024; Alvarado et al., 2024).
Note. —
@@
| (cm s-2) | 4.90.4 | 5.20.5 | 4.90.4 | |||||
@@
Note. — Best model(s) are for NIRSpec fits, and are indicated by labels E = Elf Owl (Mukherjee et al., 2024), D = Diamondback (Morley et al., 2024), and S = SAND (Gerasimov et al., 2024; Alvarado et al., 2024). When more than one model is listed, models outside of parentheses have indistinguishable best-fit , while models in parentheses have slightly higher (more poorly fit) best-fit . Model parameters are determined from an uncertainty-weighted mean of the the posterior distributions of all the best-fit models listed (see Eqns.5 and 3). The adopted distance is an uncertainty-weighted average of the model fit distance () and spectrophotometric distance () assuming the larger of the error bounds.
@@ 5.1 Population Assignments
The large distances and high Galactic latitudes of these sources translate into large vertical scaleheights from the plane, giving them a high probability of belonging to the Milky Way’s thick disk or halo populations (Burgasser et al., 2024). To quantify population membership, we used the Galactic star density model of Jurić et al. (2008) to estimate the relative space densities of thin disk, thick disk, and halo sources in the directions and at the distances for each of our targets, sampling distance uncertainties using Monte Carlo methods. The resulting probabilities are listed in Table 6. Not surprisingly, the most distant sources in our sample, RUBIES-BD-3, -4, and -5, have high probabilities of thick disk membership, with RUBIES-BD-4 being sufficiently far to have a significant probability (357%) of halo membership. This subsample includes both the classified L subdwarf RUBIES-BD-3 and the only source in the sample whose modeled parameters indicate subsolar metallicity, RUBIES-BD-5. The mid- to late-T dwarfs in our sample—RUBIES-BD-1, -6, and -7—are intrinsically fainter and detected at vertical scaleheights 1 kpc, and are thus more likely to be thin disk dwarfs, as supported by their spectral similarity to local spectral standards. The mix of thin disk, thick disk, and halo brown dwarfs mimics the composition of distant stars and brown dwarf detected in other deep JWST survey samples (Burgasser et al., 2024; Hainline et al., 2024a, b, 2025).

### v2

@@ 3.2 Markov Chain Monte Carlo Fitting
We used the ucdmcmc package (ucdmcmc) to fit the three model sets to our brown dwarf spectra. Spectra and models were compared using the same chi-square statistic as Eqn. 1. A first fit was conducted to identify the best individual model within a model set based on the minimum value. These model parameters were then used to initiate a Metropolis-Hastings Markov Chain Monte Carlo (MCMC) algorithm (1953JChPh..21.1087M; HASTINGS01041970) to determine the overall best-fit parameters and their uncertainties from the posterior distributions. Model logarithmic flux densities were linearly interpolated between grid points over logarithmic parameter values (e.g., interpolates as log ). The MCMC algorithm proposes a new set of parameters by varying the previous set using normal distributions with the fixed widths listed in Table 3. The difference between the values of the proposed () and previous () parameters were compared to the overall minimum , which in turn was compared to a random number drawn from a uniform distribution between 0 and a threshold value :
@@ 4 Results
Best fit models for each source and model set are displayed in Figures 4 and 5 and fit parameters are listed in Tables 4 and 5 for the 0.9–2.4 m (NIR) and 0.9–5.1 m (NIRSpec) regions, respectively. A summary of fit parameters across models is provided in Table 6, while a full list of parameter values and distribution visualizations are provided in a separate Zenodo link.444 Here, we describe the outcomes of the fits and the physical interpretation of the fit parameters for individual sources.
@@ 4.1 RUBIES-BD-1
The best NIR and NIRSpec fits for this T7 dwarf are from the Elf Owl and Diamondback models, with a slight preference for the former with modest vertical mixing ( = 3.11.0 averaged between NIR and NIRSpec fits). The Diamondback models require very efficient sedimentation or no clouds ( = 9.50.4). Averaging the results of the NIRSpec fits for these models yields Teff = 930 K, = 4.90.4, and [M/H] = +0.1, Elf Owl models a solar C/O ratio. The inferred effective temperature is slightly higher but consistent with previously established trends in spectral type (e.g., Teff = 830130 K for a T70.5 dwarf based on 2015ApJ...810..158F). This modest overshoot may be due to the best-fitting models failing to match the depth of the 1.6 m CH4 band or the strong CO2 band at 4.1 m (see discussion below). Nevertheless, the agreement in classification and model fit parameters indicate that this is a normal late T-type brown dwarf.
RUBIES-BD-1 was as CEERS-EGS-BD-4 in 2024ApJ...964...66H, identified from multi-band JWST/NIRCam photometry. The photometric measurements were fit to three sets of atmosphere models: Sonora Bobcat and Sonora Cholla (2021ApJ...923..269K), precursors to Diamondback and Elf Owl models; and ATMO (2020A&A...637A..38P). The fits yield Teff = 1000–1050 K and = 5.0, but no evaluation of metallicity. The authors infer a distance of 1350–1831 pc for estimated radii of 0.92–1.2 RJup, somewhat further than our estimate of 1030140 pc using a smaller radius estimate. RUBIES JWST/NIRSpec spectral data of this source were reported in 2025ApJ...980..230T, where the source is identified as o006_s00089. These authors compared the NIRSpec spectrum to Elf Owl and ATMO 2020++ (2023AJ....166...57M) model sets, the latter an updated version of the ATMO models. For the , the authors inferred a similar Teff = 955 K but a considerably smaller = 4.34, outside the range considered in these fits, albeit formally consistent with our uncertainties. also infer a much larger distance for this source, 2.00.2 kpc, by assuming a radius of 1.2 RJup as suggested by the low surface gravity, but estimate a shorter distance of 1.4–1.7 kpc if the radius is constrained to 0.8–0.9 RJup. The authors find a significantly supersolar metallicity of [M/H] = +0.410.05 from their model fits, above but again formally consistent with our fits. 2025ApJ...980..230T note that their parameter uncertainties, which are considerably smaller than those reported here, may be underestimated.
@@ 4.2 RUBIES-BD-2
The NIR spectrum for this L1 dwarf is best fit by the Diamondback models, while the full NIRSpec spectrum has equal values between the Diamondback and SAND models. As both models include a prescription for cloud condensation and scattering, it is likely the presence of clouds that rules out the condensate-free Elf Owl models as a reliable fit. Indeed, Diamondback model fits to both the NIR and NIRSpec spectra require relatively thick clouds ( = 1.81.0). However, despite providing the best fits, Figure 5 shows that the Diamondback and SAND models have difficulty reproducing the 0.9–2.4 m spectral slope of RUBIES-BD-2, the former underestimating the 1.6 and 2.1 m flux peaks and the latter overestimating the 1.2 and 1.6 m peaks. We also find that the fits between the NIR and NIRSpec ranges yield marginally distinct temperatures, with Teff = 1760 K for the former (Diamondback only) and Teff = 2040 K for the latter (average of Diamondback and SAND), a 1.7 difference. The higher temperature is more in line with local L10.5 dwarfs (Teff = 2100130 K from 2015ApJ...810..158F). We also find a 3 difference in the inferred value between the SAND models ( = 4.22) and Diamondback models ( = 5.30) based on fits to the NIRSpec data; and a 4 discrepancy between our model-scaled distance estimate of 860 pc and our spectrophotometric distance estimate of 1320100 pc. These deviations suggest that the Elf Owl, Diamondback, and SAND models all have difficulty reproducing the spectrum of RUBIES-BD-2, indicating potentially missing elements to the models.
RUBIES JWST/NIRSpec spectral data of this source were in 2025ApJ...980..230T, where the source is identified as o006_s35616. The authors compared these data to Elf Owl and BT-Settl CIFIST models (2012RSPTA.370.2765A; 2011SoPh..268..255C), finding Teff = 1810–1910 K, lower but consistent with our NIRSpec fits. However, the authors again infer a very low = 3.250.01 and a supersolar [M/H] = +1.000.01 from their Elf Owl fits, inconsistent with our findings. 2025ApJ...980..230T note significant residuals in their model fits, recognizing that neither the Elf Owl nor BT-Settl models include condensate opacity. They introduce a reddening parameter based on the interstellar extinction law of 2024ApJ...964L...3W, and find that = 3 mag substantially improves their Elf Owl model fits while yielding a substantially higher Teff = 2286 K and subsolar [M/H] = 0.710.02. It is likely that a combination of atmospheric condensate opacity and interstellar reddening are required to accurately reproduce these data.
@@ 4.3 RUBIES-BD-3
As discussed in Section 2, RUBIES-BD-3 is better matched to subdwarf standards, suggesting that it should be well-matched to metal-poor atmosphere models. However, both NIR and NIRSpec fits yield consistently better fits to supersolar metallicity models. In the NIR range, Elf Owl, Diamondback, and SAND models all yield comparable best-fit values; in the NIRSpec range, the Elf Owl models are somewhat less reliable. Primary fit parameters of Teff, , and [M/H] are in agreement across all three models and fit ranges, with a relatively high Teff = 2170130 K from the NIRSpec fits, notably warmer than the earlier-type L dwarf RUBIES-BD-2. Secondary parameters are similar to NIRSpec fits for RUBIES-BD-2, with all three models indicating relatively low surface gravities ( = 4.70.4), and for the Diamondback models thick clouds ( = 2.1). The Elf Owl model fits also indicate a high C/O = 0.73 and modest vertical mixing ( = 3.6). However, like RUBIES-BD-2, all three models have clear flaws in reproducing the observed spectrum, particularly in the NIR range where Elf Owl models show excess flux shortward of 1.3 m, and SAND models fail to match the shape of the 1.6 m peak. None of the models accurately reproduce the 1.2 m “notch” absorption feature that is characteristic of L- and T-type subdwarfs (Figure 3). Again, the models deployed for this seemingly metal-poor L dwarf may be missing a critical element of atmospheric chemistry, composition, or dynamics that may be skewing the fits toward higher metallicities.
2025ApJ...980..230T reported RUBIES JWST/NIRSpec spectral data for this source, identified as o005_s41280, and compared the data to Elf Owl and BT-Settl models. The authors find a similar Teff = 2160–2260 K from these fits without taking into account reddening, with an approximately 200 K increase when a modest reddening of = 1.0-1.5 mag is taken into account. In both cases, the authors find a very low surface gravity for this source, = 3.3–3.5, well outside our fitting parameter range. As with RUBIES-BD-2, such a low surface gravity is highly unlikely for a field brown dwarf as it would require an age 1 Myr (2023A&A...671A.119C).
@@ 4.4 RUBIES-BD-4
The relatively low S/N of the RUBIES-BD-4 spectrum results in equivalent matches among all three models in both NIR and NIRSpec ranges, and relatively large uncertainties in inferred parameters. The poorly-constrained Teff = 1800 K is higher but not inconsistent with trends among solar metallicity dwarfs for its T23 classification (Teff = 1180200 K; 2015ApJ...810..158F). Notably, SAND model fits to the NIRSpec data yield a considerably lower Teff = 1050 K as compared to the other models, even though the best-fit model has a Teff = 1506 K (Figure 5). Secondary parameters indicate efficient vertical mixing ( = 7.81.0 from Elf Owl models), thin clouds ( = 8.40.7 from Diamondback models), and near-solar elemental abundances ([M/H] = 0.1 from all models, C/O = 0.54 from Elf Owl models, and [/M] = +0.17 from SAND models). The large range in temperatures across the models results in a comparably large uncertainty in the estimated distance, from 2000 pc for SAND to 6400 pc for Diamondback, which straddle the uncertain 5400 pc distance estimate inferred from its spectral type and NIRCam photometry (Table 2).
@@ 4.5 RUBIES-BD-5
RUBIES-BD-5 is best fit in both the NIR and NIRspec ranges by the SAND models, yielding an atmospheric temperature Teff = 1270 K that is in line with its T12 classification (Teff = 1220150 K; 2015ApJ...810..158F). Figure 5 shows that the SAND models do a better job reproducing the 1–2.2 m and 4.0 m peaks compared to the Elf Owl and Diamondback models, both of which produce overly blue spectra. The NIRSpec SAND fits also yield a significantly subsolar metallicity ([M/H] = 0.7) and modest alpha enrichment ([/M] = +0.19), both indicative of a thick disk or halo origin.
@@ 4.6 RUBIES-BD-6
RUBIES-BD-6, like RUBIES-BD-4, has a relatively low S/N spectrum and is equally-well fit by all three models in both NIR and NIRSpec ranges, with a slight preference for the Elf Owl models in the latter. The NIRSpec fits yield an average Teff = 830130 K, high but consistent with its T81 classification (Teff = 680190 K; 2015ApJ...810..158F). Given the large uncertainties, there is no evidence of a metallicity or elemental compositions significantly distinct from solar, while the Diamondback models require efficient sedimentation ( = 8.0) as is typical for late T dwarfs.
@@ 4.7 RUBIES-BD-7
RUBIES-BD-7 is best fit in both the NIR and NIRSpec ranges by the Elf Owl and Diamondback models, the latter requiring very efficient sedimentation ( = 9.60.4, consistent with no clouds). The inferred Teff = 118080 K is consistent with other T50.5 dwarfs (Teff = 1030120 K; 2015ApJ...810..158F), while the metallicity ([M/H] = 0.20.3) and Elf Owl C/O = 0.48 are consistent with solar composition. The excellent agreement between the model scaling and spectrophotometric distances further indicate that this is likely a normal thin disk brown dwarf.
@@
Note. — Model labels are: E = Elf Owl (2024ApJ...963...73M), D = Diamondback (Morley_2024), and S = SAND (2024ApJ...971...65G; 2024RNAAS...8..134A).
@@
Note. — Model labels are: E = Elf Owl (2024ApJ...963...73M), D = Diamondback (Morley_2024), and S = SAND (2024ApJ...971...65G; 2024RNAAS...8..134A).
@@
| (cm s-2) | 4.90.4 | aaThe SAND NIRSpec fit included in this average has a significantly discrepant value of . | 5.20.5 | 4.90.4 | ||||
@@
Note. — Best model(s) are for NIRSpec fits, and are indicated by labels E = Elf Owl (2024ApJ...963...73M), D = Diamondback (Morley_2024), and S = SAND (2024ApJ...971...65G; 2024RNAAS...8..134A). When more than one model is listed, models outside of parentheses have indistinguishable best-fit , while models in parentheses have slightly higher (more poorly fit) best-fit . Model parameters are determined from an uncertainty-weighted mean of the posterior distributions of all the best-fit models listed (see Eqns.5 and 3). The adopted distance is an uncertainty-weighted average of the model fit distance () and spectrophotometric distance () assuming the larger of the error bounds.
@@ 5.1 Population Assignments
The large distances and high Galactic latitudes of these sources translate into large vertical scaleheights from the plane, giving them a high probability of belonging to the Milky Way’s thick disk or halo populations (2024ApJ...962..177B). To quantify population membership, we used the star density model of 2008ApJ...673..864J to estimate the relative space densities of thin disk, thick disk, and halo sources in the directions and at the distances for each of our targets, sampling distance uncertainties using Monte Carlo methods. The resulting probabilities are listed in Table 6. Not surprisingly, the most distant sources in our sample, RUBIES-BD-3, -4, and -5, have high probabilities of thick disk membership, with RUBIES-BD-4 a significant probability (357%) of halo membership. This subsample includes both the classified L subdwarf RUBIES-BD-3 and the only source in the sample whose modeled parameters indicate subsolar metallicity, RUBIES-BD-5, The mid- to late-T dwarfs in our sample—RUBIES-BD-1, -6, and -7—are intrinsically fainter and detected at vertical scaleheights 1 kpc, and are thus more likely to be thin disk dwarfs, as supported by their spectral similarity to local spectral standards. The mix of thin disk, thick disk, and halo brown dwarfs mimics the composition of distant stars and brown dwarfs detected in other deep JWST survey samples (2024ApJ...962..177B; 2024ApJ...964...66H; 2024ApJ...975...31H; 2025arXiv251000111H).